# Rediscovering Entropy Regularization: Adaptive Coefficient Unlocks Its Potential for LLM Reinforcement Learning

### v1

@@ Abstract
22footnotetext: Xiaoyun Zhang and Xiaojian Yuan contributed equally to this work. Work done during an internship at StepFun.11footnotetext: Corresponding author.33footnotetext: Under review at ACL Rolling Review, October 2025.
@@ 1 Introduction
In this work, we revisit entropy regularization in the context of RLVR for LLMs and argue that its potential has been largely underestimated due to the limitations of fixed-coefficient designs. Motivated by this concern, we conduct preliminary experiments and analysis in Section 3 and have two observations: (i) tasks of different difficulty levels require distinct exploration intensities, suggesting the need for a difficulty-aware mechanism that enables sample-level control of entropy regularization; and (ii) effective exploration during training requires maintaining the policy entropy at a specific target value below its initial entropy,
@@ 3 Preliminary Analysis
Intuitively, the degree of exploration should correlate with task difficulty: excessive exploration on easy tasks may introduce unnecessary randomness and hinder convergence, whereas difficult tasks often require stronger exploration to escape local optima and discover effective reasoning trajectories (Li et al., 2025a, b). To examine this intuition, we trained Qwen3-4B-Base with GRPO on mathematical datasets of different difficulty levels [^1] The easy task uses GSM8K, while the hard task consists of a mixture of AIME and AMC datasets., varying the strength of entropy regularization through different coefficients. As shown in Figure 2(a) and Figure 2(b), increasing the entropy coefficient improves test accuracy on the harder dataset, accompanied by a longer average token length—indicating that moderate promotion of exploration benefits challenging reasoning tasks. In contrast, on the easier dataset, stronger entropy regularization leads to a decline in accuracy, where excessive exploration with longer responses may prevent convergence toward concise and correct reasoning trajectories. These results demonstrate that the optimal level of exploration varies with task difficulty, highlighting the necessity of a difficulty-aware mechanism for entropy regularization. In addition, when the coefficient is set to an overly large value (i.e., $0.1$), both datasets showed a sudden drop in accuracy accompanied by a sharp increase in the average token length, indicating that excessive exploration may trigger entropy explosion, leading to instability in the training process.
@@ 4.1 Notations and Preliminaries
We consider reinforcement learning with verifiable rewards (RLVR) for training LLMs. Given a question-answer pair $(q,a)$ from the dataset $\mathcal{D}$, the policy model $\pi_{\theta}(\cdot\mid q)$ with parameters $\theta$ generates a response $o=(o_{1},\dots,o_{T})$. A rule-based reward function $r(q,o)\in\{0,1\}$ provides binary correctness judgment.
@@
$\hat{A}_{i,t}=\frac{r(q,o_{i})-\text{mean}(\{R_{i}\}_{i=1}^{G})}{\text{std}(\{R_{i}\}_{i=1}^{G})}.$ (1)
@@
${\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left(\operatorname{min}\left(w_{i,t}(\theta)\hat{A}_{i,t},\right.\right.\right.\text{ clip }}$
$\left.\left.\left.\left(w_{i,t}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i,t}\right)-\beta D_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\text{ref }}\right)\right)\right]$
where $w_{i,t}(\theta)=\frac{\pi_{\theta}\left(o_{i,t}\mid q,o_{i,<t}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,t}\mid q,o_{i,<t}\right)}$, $\pi_{\text{ref}}$ denotes the reference policy and $\beta$ controls the strength of the KL penalty.
@@
To alleviate the policy entropy collapse, a common approach is to explicitly incorporate entropy regularization as an auxiliary objective. Formally, for an autoregressive policy $\pi_{\theta}$, the token-level entropy at step $t$ is defined as
$\mathcal{H}_{t}=-\!\!\sum_{a\in\mathcal{V}}\pi_{\theta}(a\mid x,y_{<t})\,\log\pi_{\theta}(a\mid x,y_{<t}),$ (3)
@@
$\mathcal{H}(\pi_{\theta}(\cdot\mid x))=\frac{1}{T}\sum_{t=1}^{T}\mathcal{H}_{t},$ (4)
with $T$ being the response length. The entropy regularization objective is given by
$\mathcal{J}_{\text{ent}}(\theta)=\gamma\,\mathbb{E}_{(x,y)\sim\mathcal{D}}\!\left[\mathcal{H}(\pi_{\theta}(\cdot\mid x))\right],$ (5)
@@
Finally, the overall training objective combines this entropy term with the policy optimization objective (e.g., GRPO):
@@ 4.2 Adaptive Entropy Regularization (AER)
$+\;\mathbb{E}_{(x,y)\sim\mathcal{D}}\!\big[\lambda_{t}(x,y)\,\mathcal{H}(\pi_{\theta}(\cdot\mid x))\big],$
where $\mathcal{H}(\pi_{\theta}(\cdot\mid q))$ denotes the sequence-level entropy and $\lambda_{t}(q,o)$ represents the adaptive coefficient for each sample.
@@
$\lambda_{t}(q,o_{j})$ $=\alpha_{t}\cdot\frac{\max\{0,\,\rho-g(q)\}}{\rho+\varepsilon}$ (9)
@@
where $\alpha_{t}$ is a global entropy scaling factor controlling the overall strength of entropy regularization, which is adaptively updated through the dynamic adjustment mechanism introduced in C3. The coefficient $\lambda_{t}(q,o_{j})$ is positive only for hard questions ($g(q)\leq\rho$) and zero otherwise. Additionally, it allocates larger entropy coefficients to more difficult questions based on the difference between the current $g(q)$ and $\rho$ to encourage more exploration.
@@
As analyzed in Section 3, maintaining the policy entropy within a certain “sweet spot” below its initial value can appropriately promote exploration and stabilize training. However, the initial policy entropy $\widehat{\mathcal{H}}_{0}$ varies substantially across base models, datasets, and sampling temperature, making it difficult to directly specify a specific value as the target entropy. Therefore, we seek an adaptive mechanism that can determine the value of the target entropy based on the initial entropy at each run.
@@
where $\tau$ is a predefined reduction ratio representing the desired relative decrease from the initial entropy. By anchoring the target entropy to the initial entropy, it adaptively calibrates the effective “exploration budget” across different setups, thereby eliminating the need for repeated hyperparameter tuning and improving cross-run stability.
@@
Definition. Let $\widehat{\mathcal{H}}_{t}$ denote the batch-level policy entropy at iteration $t$. With step size $\eta>0$, the global entropy scaling factor $\alpha_{t}$ in Equation 9 is updated as
@@
Each training iteration of AER proceeds as follows: (i) estimate the group accuracy $g(q)$ for each question via Equation 8; (ii) compute difficulty-aware coefficients $\lambda_{t}(q,o)$ via  Equation 9; (iii) optimize $\mathcal{J}_{\text{AER}}(\theta)$ in  Equation 7; and (iv) update $\alpha_{t}$ according to  Equation 11. This closed-loop process adaptively allocates and regulates exploration, maintaining stable entropy dynamics throughout RLVR training.
@@ 5.1 Experiment Setup
We compare AER against several baselines, including the standard GRPO (Shao et al., 2024), GRPO with the Clip-Higher (Yu et al., 2025) and several advanced exploration-oriented methods: Ent-Adv (Cheng et al., 2025), which introduces an entropy-based advantage term to encourage longer reasoning trajectories, and Clip-Cov and KL-Cov (Cui et al., 2025b), which regularize policy updates through token–entropy covariance to stabilize training.
@@ 5.2 Main Results
Following prior work (Yue et al., 2025; Cheng et al., 2025; Jiang et al., 2025b), we further examine the exploration boundary of different methods through pass@$k$ performance, which reflects the model’s upper bound of reasoning capability. Specifically, we evaluate the Qwen3-8B-Base model on two challenging benchmarks—AIME2024 and AIME2025—extending $k$ up to $256$ to assess how performance scales with increased rollout attempts. As shown in Figure 3, AER consistently achieves higher pass@$k$ performance across all $k$, indicating stronger exploration and reasoning diversity. The advantage becomes more pronounced as $k$ increases on the AIME25 dataset, where AER maintains a clear margin over baselines throughout the curve. This trend suggests that AER more effectively broadens the search space of reasoning trajectories, enabling the policy to discover a larger proportion of correct solutions under extended sampling. Detailed quantitative results can be found in Appendix C.1.
@@
In Figure 3(b), the training reward of AER reaches a higher level than the baselines, suggesting that our method facilitates more effective policy improvement during training. As shown in Figure 3(c), the policy entropy of naive GRPO drops sharply within the first 100 steps, indicating early entropy collapse. Although the baseline slightly mitigates this effect, its entropy still gradually declines over steps. AER maintains the policy entropy at a moderate target level throughout training, which help sustain consistent and balanced exploration. Despite maintaining higher entropy, AER produces shorter average response lengths than the baseline as shown in Figure 3(e), implying that our method guides exploration more effectively without unnecessary verbosity. Finally, Figure 3(f) shows that AER dynamically maintains a stable policy entropy throughout training and allocates more exploration to difficult samples, thereby facilitating the policy to escape local optima and continuously improve performance. Overall, these results suggest that AER promotes stable entropy dynamics, balanced exploration, and sustained performance gains during RLVR training.
@@ 5.3 Ablation Study
Fixing $\rho{=}0.2$, we sweep $\tau$ to study how anchoring the target entropy to the initial entropy affects performance (Figure 4(b)). The Avg. pass@1 curve exhibits a clear sweet spot at $\tau{=}0.4$ (55.4), with a mild plateau for $\tau\!\in\![0.5,0.6]$ and degradation in both directions: under-reduction at $\tau\!\leq\!0.3$ and over-preservation at high-entropy settings $\tau\!\geq\!0.8$. These results substantiate the empirical premise that moderate entropy reduction accompanies accuracy gains: compared with naïve GRPO, AER stabilizes the global entropy near the prescribed band, retaining sufficient diversity while avoiding premature collapse. Pushing $\tau$ too high inflates entropy beyond the useful range, diluting the signal for exploitation; pushing it too low curtails beneficial exploration. In practice, we thus adopt $\tau\!\in\![0.4,0.6]$ as default.
@@
To test whether AER’s closed-loop entropy control transfers across data and architectures without retuning, we replace DeepscaleR with dapo_math_17k and train Qwen3-8B-Base and Qwen2.5-7B-Base using the same default hyperparameters ($\rho{=}0.2,\ \tau{=}0.4$). The test-score trajectories in Figure 4(c) and Figure 4(d) show that AER achieves faster early gains and maintains a clear, persistent margin over Clip-Higher throughout training, while converging to a higher (or comparable) asymptote than GRPO. This behavior indicates that the initial-anchored target entropy (C2) together with dynamic global scaling (C3) adapts the effective exploration level to the new data distribution and model backbone.
@@ A.2 Hyperparameters
w/AER 128 32 0.0 8k 1e-6 30 20 $t{=}0.6,\beta{=}0.2,\eta{=}0.005$
Table 3: Training configurations. We keep the compute budget and decoding policy matched across methods. The KL coefficient controls regularization strength in PPO-style objectives; DAPO removes KL entirely. Clip-Higher uses asymmetric clipping ratios ($c_{l}$, $c_{h}$). Ent-Adv introduces entropy-based advantage estimation with scaling factor $\alpha$ and temperature $\kappa$. KL-Cov and Clip-Cov add covariance regularization with coefficients $\rho_{kl}$ and $\rho_{clip}$ respectively. AER (Adaptive Entropy Regulation) introduces adaptive entropy control with target entropy $t$, difficulty threshold $\beta$, and adaptation rate $\eta$.
@@
  • AER: Implements adaptive entropy regulation with target entropy $t=0.6$, difficulty threshold $\beta=0.2$ for distinguishing hard vs. easy samples, and adaptation rate $\eta=0.005$ for dynamic parameter adjustment. The method automatically maintains the target entropy level through adaptive $\alpha$ parameter updates.
@@
All ablations modify a single factor at a time (e.g., toggling KL, changing $t$ in APARL) while keeping the data curriculum, tokenizer, context length, and decoding policy fixed. This controls confounders and isolates the effect of exploration–exploitation regularization on pass@ $1$ and pass@ $k$.

### v2

@@ Abstract
22footnotetext: Xiaoyun Zhang and Xiaojian Yuan contributed equally to this work. Work done during an internship at StepFun.11footnotetext: Corresponding author.
@@ 1 Introduction
In this work, we revisit entropy regularization in the context of RLVR for LLMs and argue that its potential has been largely underestimated due to the limitations of fixed-coefficient designs. Motivated by this concern, we conduct preliminary analysis in Section 3 and have two observations: (i) tasks of different difficulty levels require distinct exploration intensities, suggesting the need for a difficulty-aware mechanism that enables sample-level control of entropy regularization; and (ii) effective exploration during training requires maintaining the policy entropy at a specific target value below its initial entropy.
@@ 3 Preliminary Analysis
Intuitively, the degree of exploration should correlate with task difficulty: excessive exploration on easy tasks may introduce unnecessary randomness and hinder convergence, whereas difficult tasks often require stronger exploration to escape local optima and discover effective reasoning trajectories (Li et al., 2025a, b). To examine this intuition, we train Qwen3-4B-Base with GRPO on mathematical datasets of different difficulty levels [^1] The easy task uses GSM8K, while the hard task consists of a mixture of AIME and AMC datasets., varying the strength of entropy regularization through different coefficients. As shown in Figure 2(a) and Figure 2(b), increasing the entropy coefficient improves test accuracy on the harder dataset, accompanied by a longer average token length—indicating that moderate promotion of exploration benefits challenging reasoning tasks. In contrast, on the easier dataset, stronger entropy regularization leads to a decline in accuracy, where excessive exploration with longer responses may prevent convergence toward concise and correct reasoning trajectories. These results demonstrate that the optimal level of exploration varies with task difficulty, highlighting the necessity of a difficulty-aware mechanism for entropy regularization. In addition, when the coefficient is set to an overly large value (i.e., $0.1$), both datasets showed a sudden drop in accuracy accompanied by a sharp increase in the average token length, indicating that excessive exploration may trigger entropy explosion, leading to instability in the training process.
@@ 4.1 Notations and Preliminaries
We consider reinforcement learning with verifiable rewards (RLVR) for training LLMs. Given a question-answer pair $(q,a)$ from the dataset $\mathcal{D}$, the policy model $\pi_{\theta}(\cdot\mid q)$ with parameters $\theta$ generates a response $o$. A rule-based reward function $r(q,o)\in\{0,1\}$ provides binary correctness judgment based on $(q,a).$
@@
$\hat{A}_{i}=\frac{r(q,o_{i})-\text{mean}(\{R_{i}\}_{i=1}^{G})}{\text{std}(\{R_{i}\}_{i=1}^{G})}.$ (1)
@@
${\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{l=1}^{|o_{i}|}\left(\operatorname{min}\left(w_{i,l}(\theta)\hat{A}_{i},\right.\right.\right.\text{ clip }}$
$\left.\left.\left.\left(w_{i,l}(\theta),1-\varepsilon,1+\varepsilon\right)\hat{A}_{i}\right)-\beta D_{\mathrm{KL}}\left(\pi_{\theta}\|\pi_{\text{ref }}\right)\right)\right]$
where $w_{i,l}(\theta)=\frac{\pi_{\theta}\left(o_{i,l}\mid q,o_{i,<l}\right)}{\pi_{\theta_{\text{old }}}\left(o_{i,l}\mid q,o_{i,<l}\right)}$, $\pi_{\text{ref}}$ denotes the reference policy and $\beta$ controls the strength of the KL penalty.
@@
To alleviate the policy entropy collapse, a common approach is to explicitly incorporate entropy regularization as an auxiliary objective. Formally, for an autoregressive policy $\pi_{\theta}$, the token-level entropy at position $l$ is defined as
$\mathcal{H}_{l}=-\!\!\sum_{o_{l}\in\mathcal{V}}\pi_{\theta}(o_{l}\mid q,o_{<l})\,\log\pi_{\theta}(o_{l}\mid q,o_{<l}),$ (3)
@@
$\mathcal{H}(\pi_{\theta}(\cdot\mid q))=\frac{1}{L}\sum_{l=1}^{L}\mathcal{H}_{l},$ (4)
with $L$ being the response length. The entropy regularization objective is given by
$\mathcal{J}_{\text{ent}}(\theta)=\gamma\,\mathbb{E}_{q\sim\mathcal{D}}\!\left[\mathcal{H}(\pi_{\theta}(\cdot\mid q))\right],$ (5)
@@
Finally, the overall training objective combines this entropy regularization with the policy optimization objective (e.g., GRPO):
@@ 4.2 Adaptive Entropy Regularization (AER)
$+\;\mathbb{E}_{q\sim\mathcal{D}}\!\big[\lambda_{t}(q)\,\mathcal{H}(\pi_{\theta}(\cdot\mid q))\big],$
where $\mathcal{H}(\pi_{\theta}(\cdot\mid q))$ denotes the sequence-level entropy and $\lambda_{t}(q)$ represents the adaptive coefficient for each sample.
@@
$\lambda_{t}(q)$ $=\alpha_{t}\cdot\frac{\max\{0,\,\rho-g(q)\}}{\rho+\varepsilon}$ (9)
@@
where $\alpha_{t}$ is a global entropy scaling factor controlling the overall strength of entropy regularization, which is adaptively updated through the dynamic adjustment mechanism introduced in C3. The coefficient $\lambda_{t}(q)$ is positive only for hard questions ($g(q)\leq\rho$) and zero otherwise. Additionally, it allocates larger entropy coefficients to more difficult questions based on the difference between the current $g(q)$ and $\rho$ to encourage more exploration.
@@
As analyzed in Section 3, maintaining the policy entropy within a certain “sweet spot” below its initial value can appropriately promote exploration and stabilize training. However, the initial policy entropy varies substantially across base models, datasets, and sampling temperature, making it difficult to directly specify a specific value as the target entropy. Therefore, we seek an adaptive mechanism that can determine the value of the target entropy based on the initial entropy at each run.
@@
where $\tau$ is a predefined reduction ratio representing the desired relative decrease from the initial entropy. By anchoring the target entropy to the initial entropy, it adaptively calibrates the effective “exploration budget” across different setups, thereby alleviating the burden of repeated hyperparameter tuning and improving cross-run stability.
@@
Definition. Let $\widehat{\mathcal{H}}_{t}$ denote the policy entropy at iteration $t$. With step size $\eta>0$, the global entropy scaling factor $\alpha_{t}$ in Equation 9 is updated as
@@
Each training iteration of AER proceeds as follows: (i) estimate the group accuracy $g(q)$ for each question via Equation 8; (ii) compute difficulty-aware coefficients $\lambda_{t}(q)$ via  Equation 9; (iii) optimize $\mathcal{J}_{\text{AER}}(\theta)$ in  Equation 7; and (iv) update $\alpha_{t}$ according to  Equation 11. This closed-loop process adaptively allocates and regulates exploration, maintaining stable entropy dynamics throughout RLVR training.
@@ 5.1 Experiment Setup
We compare AER against several baselines, including the vanilla GRPO (Shao et al., 2024), GRPO with Clip-Higher (Yu et al., 2025) and several advanced exploration-oriented methods: Ent-Adv (Cheng et al., 2025), which introduces an entropy-based advantage term to encourage longer reasoning trajectories, and Clip-Cov and KL-Cov (Cui et al., 2025b), which regularize policy updates through token–entropy covariance to stabilize training.
@@ 5.2 Main Results
Following prior work (Yue et al., 2025; Cheng et al., 2025; Jiang et al., 2025b), we further examine the exploration boundary of different methods through pass@$k$ performance, which reflects the model’s upper bound of reasoning capability. Specifically, we evaluate the Qwen3-8B-Base model on two challenging benchmarks—AIME2024 and AIME2025—extending $k$ up to $256$ to assess how performance scales with increased rollout attempts. As shown in Figure 3(a) and Figure 3(d), AER consistently achieves higher pass@$k$ performance across all $k$, indicating stronger exploration and reasoning diversity. The advantage becomes more pronounced as $k$ increases on the AIME25 dataset, where AER maintains a clear margin over baselines throughout the curve. This trend suggests that AER more effectively broadens the search space of reasoning trajectories, enabling the policy to discover a larger proportion of correct solutions under extended sampling. Detailed quantitative results can be found in Appendix C.1.
@@
In Figure 3(b), the training reward of AER reaches a higher level than the baselines, suggesting that our method facilitates more effective policy improvement during training. As shown in Figure 3(c), the policy entropy of vanilla GRPO drops sharply within the first 100 steps, indicating early entropy collapse. Although the baseline slightly mitigates this effect, its entropy still gradually declines over steps. AER maintains the policy entropy at a moderate target level throughout training, which help sustain consistent and balanced exploration. Despite maintaining higher entropy, AER produces shorter average response lengths than the baseline as shown in Figure 3(e), implying that our method guides exploration more effectively without unnecessary verbosity. Finally, Figure 3(f) shows that AER dynamically maintains a stable policy entropy throughout training and allocates more exploration to difficult samples, thereby facilitating the policy to escape local optima and continuously improve performance. Overall, these results suggest that AER promotes stable entropy dynamics, balanced exploration, and sustained performance gains during RLVR training.
@@ 5.3 Ablation Study
Fixing $\rho{=}0.2$, we sweep $\tau$ to study how anchoring the target entropy to the initial entropy affects performance (Figure 4(b)). The Avg. pass@1 curve exhibits a clear sweet spot at $\tau{=}0.4$ (55.4), with a mild plateau for $\tau\!\in\![0.5,0.6]$ and degradation in both directions: under-reduction at $\tau\!\leq\!0.3$ and over-preservation at high-entropy settings $\tau\!\geq\!0.8$. These results substantiate the empirical premise that moderate entropy reduction accompanies accuracy gains: compared with vanilla GRPO, AER stabilizes the global entropy near the prescribed band, retaining sufficient diversity while avoiding premature collapse. Pushing $\tau$ too high inflates entropy beyond the useful range, diluting the signal for exploitation; pushing it too low curtails beneficial exploration. In practice, we thus adopt $\tau\!\in\![0.4,0.6]$ as default.
@@
To test whether AER’s closed-loop entropy control transfers across data and architectures without retuning, we replace DeepscaleR with dapo_math_17k and train Qwen3-8B-Base and Qwen2.5-7B-Base using the same default hyperparameters ($\rho{=}0.2,\ \tau{=}0.4$). The test-score curves in Figure 4(c) and Figure 4(d) show that AER achieves faster early gains and maintains a clear, persistent margin over Clip-Higher throughout training, while converging to a higher (or comparable) asymptote than GRPO. This behavior indicates that the initial-anchored target entropy (C2) together with dynamic global scaling (C3) adapts the effective exploration level to the new data distribution and model backbone.
@@ A.2 Hyperparameters
w/AER 128 32 0.0 8k 1e-6 30 20 $\tau{=}0.4,\rho{=}0.2,\eta{=}0.005$
Table 3: Training configurations. We keep the compute budget and decoding policy matched across methods. The KL coefficient controls regularization strength in PPO-style objectives; DAPO removes KL entirely. Clip-Higher uses asymmetric clipping ratios ($c_{l}$, $c_{h}$). Ent-Adv introduces entropy-based advantage estimation with scaling factor $\alpha$ and temperature $\kappa$. KL-Cov and Clip-Cov add covariance regularization with coefficients $\rho_{kl}$ and $\rho_{clip}$ respectively. AER (Adaptive Entropy Regulation) introduces adaptive entropy control with target entropy $\tau$, difficulty threshold $\rho$, and adaptation rate $\eta$.
@@
  • AER: Implements adaptive entropy regulation with $\tau=0.4$, difficulty threshold $\rho=0.2$ for distinguishing hard vs. easy samples, and adaptation rate $\eta=0.005$ for dynamic parameter adjustment. The method automatically maintains the target entropy level through adaptive $\alpha$ parameter updates.
@@
All ablations modify a single factor at a time (e.g., toggling KL, changing $\tau$ in AER) while keeping the data curriculum, tokenizer, context length, and decoding policy fixed. This controls confounders and isolates the effect of exploration–exploitation regularization on pass@ $1$ and pass@ $k$.